Given that z = ax + √(a2 − 4y) + c is the complete integral of the PDE, p2-q2 = 4 ,determine its general integral.


1
Expert's answer
2021-03-09T17:15:47-0500

Given that "z=ax+\sqrt{a^2-4y}+c" is the complete integral of the PDE, "p^2-q^2=4"

To find the general solution we put "c=f(a)," where "f" is an arbitrary function



"\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ z=ax+\sqrt{a^2-4y}+f(a) \ \ \ \ \ \ \ \ \ \ \ \ \ (1)"

Differentiate "(1)" partially with respect to "a"



"x+{2a \over 2\sqrt{a^2-4y}}+f'(a)=0""{a \over \sqrt{a^2-4y}}=-(x+f'(a))""a^2=a^2(x+f'(a))^2-4y(x+f'(a))^2""a=\pm2\sqrt{{(x+f'(a))^2y \over (x+f'(a))^2-1}}""z=ax-{a \over x+f'(a)}+f(a)"

The general solution is



"z=\pm2\sqrt{{(x+f'(a))^2y \over (x+f'(a))^2-1}}x\mp\text{sgn}(x+f'(a))2\sqrt{{y \over (x+f'(a))^2-1}}+f(a)"